OpenStudy (anonymous):

What is the value of x? -2/7x +6 +1/7x =18

OpenStudy (michele_laino):

is this your equation? \[-\frac{ 2 }{ 7 }x +6 +\frac{ 1 }{ 7 }x=18\]

OpenStudy (anonymous):

Yes

OpenStudy (michele_laino):

Ok! Now, please multiply by 7 both sides of your equation, then write what you get

OpenStudy (anonymous):

-2/7x +6x = 126

OpenStudy (michele_laino):

I think not, because you have to performthe subsequent calculus: \[7*(-\frac{ 2 }{ 7 }x+6+\frac{ 1 }{ 7 }x)=7*18\]

OpenStudy (michele_laino):

for example: \[7*(-\frac{ 2 }{ 7 }x)=-2x\] now try, pleae

OpenStudy (anonymous):

Do you have to divide both sides by 7?

OpenStudy (michele_laino):

No, I have to multiply both sides by 7, as I wrote above

OpenStudy (anonymous):

Okay, so it's -2/7x = -14x @Michele_Laino

OpenStudy (michele_laino):

\[-2*x+42+x=126\] is what you get after multiplication by 7

OpenStudy (anonymous):

I don't understand any of this..

OpenStudy (michele_laino):

don't warry, your answer is: x=-84

OpenStudy (michele_laino):

oops, don't worry....x=-84

OpenStudy (anonymous):

Thank you for trying to help me!

OpenStudy (michele_laino):

thank you!